295 gram Rubidium Oxide reacted with 202 gram calcium cyanide to produce Rubidium cyanide  and Calcium  Oxide.  Find

a)gram  of  Calcium  oxide

b) gram of excess  reactant  used up

c) gram of  excess  reactant that remained  unreacted




1
Expert's answer
2021-09-07T23:28:07-0400

Solution:

Calculate moles of each reactant:

(295 g Rb2O) × (1 mol Rb2O / 186.94 g Rb2O) = 1.578 mol Rb2O

(202 g Ca(CN)2) × (1 mol Ca(CN)2 / 92.1128 g Ca(CN)2) = 2.193 mol Ca(CN)2


Balanced chemical equation:

Rb2O + Ca(CN)2 → 2RbCN + CaO


According to stoichiometry:

1 mol of Rb2O reacts with 1 mol of Ca(CN)2.

Thus, 1.578 mol of Rb2O reacts with:

(1.578 mol Rb2O) × (1 mol Ca(CN)2 / 1 mol Rb2O) = 1.578 mol Ca(CN)2

However, initially there is 2.193 mol of Ca(CN)2 (according to the task).

Therefore, Rb2O acts as limiting reagent and Ca(CN)2 is excess reactant.


According to stoichiometry:

1 mol of Rb2O produces 1 mol of CaO.

Thus, 1.578 mol of Rb2O produces:

(1.578 mol Rb2O) × (1 mol CaO / 1 mol Rb2O) = 1.578 mol CaO


Calculate the mass of calcium oxide (CaO):

(1.578 mol CaO) × (56.0774 g CaO / 1 mol CaO) = 88.49 g CaO = 88.5 g CaO


Calculate the mass of excess reactant used up (Ca(CN)2):

(1.578 mol Ca(CN)2) × (92.1128 g Ca(CN)2 / 1 mol Ca(CN)2) = 145.35 g Ca(CN)2


Calculate the mass of excess reactant that remained unreacted (Ca(CN)2):

202 g Ca(CN)2 - 145.35 g Ca(CN)2 = 56.65 g Ca(CN)2


Answers:

a) 88.5 g CaO

b) 145.35 g Ca(CN)2

c) 56.65 g Ca(CN)2